# No Size Fits All: The Perils and Pitfalls of Leveraging LLMs Vary with Company Size

### v1

@@ 8 Ethics Statement
To our knowledge, this study presents minimal ethical concerns. However, to maintain transparency, we provide a detailed analysis of all 68 papers present in the survey in Section C. Each paper is reviewed by at least three individuals to validate its claims and findings. We conduct the industrial case study, following the guidelines outlined by the ACL ethics review policy  [^2] https://aclrollingreview.org/ethicsreviewertutorial, thereby Ethics Review Boards (ERB) approval is not necessary. It’s important to note that our research involving human subjects does not entail the collection of any medical or sensitive information from the users.
@@ Appendix B Industrial Case Study Details
1. Participant level of expertise in LLMs? $\square$ Beginner $\square$ Intermediate $\square$ Proficient $\square$ Expert $\square$ NA 2. Application Domain $\square$ Healthcare $\square$ Banking $\square$ Financial $\square$ Retail $\square$ Security $\square$ Privacy $\square$ Legal $\square$ Marketing & Advertising $\square$ Education $\square$ Media and entertainment $\square$ Human Resources(HR) $\square$ eCommerce $\square$ Other:   3. What is the name of task that LLM(s) performs in your project? 4. Type of data used? $\square$ Tabular $\square$ Image $\square$ Video $\square$ Audio $\square$ Text $\square$ More than one modality $\square$ Other:   5. How are the LLMs used? $\square$ Fine-tuning $\square$ Zero-shot or Few-shot $\square$ In-context learning $\square$ Other:
@@ Appendix C Survey Papers Checklist
38 Liu et al. (2023c) CHI 2023 GPT3 Google Research Generic Retrieval

### v2

@@ 8 Ethics Statement
To our knowledge, this study presents minimal ethical concerns. However, to maintain transparency, we provide a detailed analysis of all 68 papers present in the survey in Appendix Section C. Each paper is reviewed by at least three individuals to validate its claims and findings. We conduct the industrial case study, following the guidelines outlined by the ACL ethics review policy  [^3] https://aclrollingreview.org/ethicsreviewertutorial, thereby Ethics Review Boards (ERB) approval is not necessary. It’s important to note that our research involving human subjects does not entail the collection of any medical or sensitive information from the users.
@@ Appendix B Industrial Case Study Details
1. Participant level of expertise in LLMs? $\square$ Beginner $\square$ Intermediate $\square$ Proficient $\square$ Expert $\square$ NA 2. Application Domain $\square$ Healthcare $\square$ Banking $\square$ Financial $\square$ Retail $\square$ Security $\square$ Privacy $\square$ Legal $\square$ Marketing & Advertising $\square$ Education $\square$ Media and entertainment $\square$ Human Resources(HR) $\square$ eCommerce $\square$ Other:   3. What is the name of the task that LLM(s) performs in your project? 4. Type of data used? $\square$ Tabular $\square$ Image $\square$ Video $\square$ Audio $\square$ Text $\square$ More than one modality $\square$ Other:   5. How are the LLMs used? $\square$ Fine-tuning $\square$ Zero-shot $\square$ In-context learning $\square$ Other:
@@ Appendix C Survey Papers Checklist
38 Liu et al. (2023b) CHI 2023 GPT3 Google Research Generic Retrieval